OpenStudy (anonymous):

It has been over 1 hour, so here is the proof:

OpenStudy (anonymous):

Find all Polynomials such that:\[f(x^2) + f(x)f(x+1) = 0\]

OpenStudy (anonymous):

therefore: \[f(x)f(x+1) = -f(x^2)\]

OpenStudy (anonymous):

substitute in: \[g(x) = -f(x)\]

OpenStudy (anonymous):

\[g(x)g(x+1) = g(x^2)\]

OpenStudy (anonymous):

now, we know that if: \[g(a) = 0 \] \[g(a^2) = 0\]

OpenStudy (anonymous):

therefore, if the modulus of a is greater than one, or between one and 0, we will have an infinite number of roots

OpenStudy (anonymous):

now, we write the relationship: \[g(x) = \frac{g(x^2)}{g(x+1)} = \frac{g((x-1)^2)}{g(x-1)}\]

OpenStudy (anonymous):

any questions so far?

OpenStudy (anonymous):

so, if a is a root, (a-1)^2 is a root

OpenStudy (anonymous):

therefore: \[|(a-1)^2| = 1,0\]

OpenStudy (anonymous):

therefore: \[|a-1|^2 = 1,0\]

OpenStudy (anonymous):

therefore: \[|a-1| = 1,0\]

OpenStudy (anonymous):

so, there are only four numbers such that: \[|a| = |a-1| = 1,0\] 1,0,e^(i*pi/3), and e^(i*5*pi/3)

OpenStudy (anonymous):